# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

### v1

@@ 2.2.1 Momentum space OPE
Figure 4: The coefficient functions for various operators at the order ${\cal O}(z^{4})$. The dashed lines are massless propagators. The big black circles are insertion of the bare mass $-m_{0}^{2}$. The double lines denote the “external soft fields” from condensates attached to the coefficient functions. The tree-level coefficient $-\frac{3z^{4}}{24d(d+2)}$ for $\phi\partial^{4}\phi$ are omitted in the figure, but crucial for coordinate space expansion.
@@
$C_{0}\left(\frac{p^{2}}{\mu^{2}},d\right)=\frac{1}{p^{2}}\left(\frac{\mu_{0}^{2}}{p^{2}}\right)^{2-d}\frac{8^{1-d}\pi^{\frac{3}{2}-d}\Gamma\left(\frac{d}{2}-1\right)^{2}}{\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{3d}{2}-3\right)\sin\pi d}\ ,$ (40)
$C_{1}\left(\frac{p^{2}}{\mu^{2}},d\right)=\frac{1}{p^{2}}\left(\frac{\mu_{0}^{2}}{p^{2}}\right)^{1-\frac{d}{2}}\frac{2^{3-2d}\pi^{\frac{1}{2}-\frac{d}{2}}\Gamma\left(2-\frac{d}{2}\right)\Gamma\left(\frac{d}{2}-1\right)}{\Gamma\left(\frac{d-1}{2}\right)}\ ,$ (41)
@@
One can show that
@@
which is exactly the expansion of $\Sigma(p^{2})$. Here we would like to interpret the above in the form of renormalized OPE. Clearly, the second term should be regarded as the OPE for $\langle\phi^{2}\rangle_{0}$, as can be easily checked. However, the last term is not all of the $\langle\phi^{4}\rangle_{0}=3I_{0}^{2}$ contribution. The reason is that there is also a term corresponding to the double mass insertion. Namely, one has $\frac{1}{4}I_{0}^{2}+\frac{1}{2}I_{0}^{2}=\frac{1}{4}\times 3I_{0}^{2}$. The above contribution is $\frac{1}{6}\times 3I_{0}^{2}$, while there is another $\frac{1}{12}\times 3I_{0}^{2}$ that should cancel with mass insertions. To include all contributions one needs the bare mass
$m_{0}^{2}=m^{2}-\frac{g}{2}I_{0}\left(\frac{\mu^{2}}{m^{2}},d\right)\ .$ (45)
It is the bare mass that enter the OPE calculation. One introduces $m_{0}^{2}=m_{r}^{2}-\frac{g}{8\pi\epsilon}$, where $m_{r}$ is the short distance mass. We would like then to express the OPE in terms of $m_{r}$ and renormalized operators. One first write in terms of bare operators
$S(p^{2})(p^{2})^{2}=\left(\frac{g^{2}}{6}C_{0}+\frac{m_{0}^{4}}{p^{2}}\right)+\left(\frac{g^{2}}{2}C_{1}+\frac{gm_{0}^{2}}{p^{2}}\right)I_{0}+\left(\frac{g^{2}}{6p^{2}}+\frac{g^{2}}{12p^{2}}\right)3I_{0}^{2}\ .$ (46)
Now, one would like to introduce renormalized operators recursively.
One introduce renormalized operators in the $\overline{MS}$ scheme
$I_{0}=\langle\phi^{2}\rangle_{r}+\frac{1}{4\pi\epsilon}\ ,$ (47)
$3I_{0}^{2}=\langle\phi^{4}\rangle_{r}+\frac{3}{2\pi\epsilon}\langle\phi^{2}\rangle_{r}+\frac{3}{(4\pi)^{2}\epsilon^{2}}\ .$ (48)
Notice that $\langle\phi^{4}\rangle_{r}=3\langle\phi^{2}\rangle_{r}^{2}$ at this order, despite possible $\frac{1}{\epsilon}\times\epsilon$ effects. The point is that $\langle\phi^{2}\rangle_{r}$ and $\langle\phi^{4}\rangle_{r}$ are finite in the $\epsilon\rightarrow 0$ limit, but in the $\epsilon\rightarrow 0$ limit they can still contain ${\cal O}(\epsilon)$ terms.
After all, one has
@@
In the $\epsilon\rightarrow 0$ limit, all of above are finite. In fact, requiring that the $\epsilon\rightarrow 0$ limit exist uniquely determines all the $\frac{1}{\epsilon}$ poles in the mixing, once the tree-level mixing constant $Z_{10}^{(0)}=\frac{1}{4\pi\epsilon}$ for $\phi^{2}$ is determined from the coordinate space expansion at $z^{0}$ order, and the one-loop level mixing $m_{0}^{2}=m_{r}^{2}-\frac{g}{8\pi\epsilon}$ is fixed from the $z^{2}$ order consideration. In particular
@@
Tp reproduce the full result, notice in the current case one has $m_{r}^{2}=m^{2}-\frac{g}{2}\langle\phi^{2}\rangle_{r}=m^{2}-\frac{g}{8\pi}\ln\frac{\mu^{2}}{m^{2}}$, and the operator condensates read
@@
Combining all above, one has the $\mu$-independent result $S_{2}(p^{2})(p^{2})^{2}=\frac{g^{2}}{32\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{m^{2}}+\frac{m^{4}}{p^{2}}$.
Here we should stress that if one only keeps the naive perturbative result $\frac{g^{2}}{6}C_{0}$ in $C_{0,r}$ without introducing additive counter-terms, then not only the $\frac{1}{\epsilon}$ poles no longer cancel, but the double log term in the finite part would also have the wrong coefficient $\frac{g^{2}}{16\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{\mu^{2}}$.
@@
It is clear now that all the $\frac{1}{\epsilon}$ poles cancel, the $\frac{1}{\epsilon^{2}}$ poles add to cancel the remaining $\frac{g^{2}}{2p^{2}}\frac{1}{(4\pi)^{2}\epsilon^{2}}$. The $\zeta_{2}$ term all cancel and the double-log terms combine to reproduce the full result for $C_{0,r}$. Therefore, in the DR calculation, it is important to keep the $d$ dependency of the bare coefficient functions to the end, and the $\frac{1}{\epsilon}$ effects can be crucial.
@@
Here we note that the above is the general form of the large $p^{2}$ expansion up to second order in $\frac{1}{p^{2}}$ for a generic $\phi_{2}^{4}$, where $\langle:\phi^{2}:\rangle$ and $\langle:\phi^{4}:\rangle$ can be non-perturbative. At small $g$, condensates for these two Wick-polynomials are of order ${\cal O}(\frac{g^{2}}{m^{4}})$. The double logarithmic dependency is always purely determined perturbatively.
@@ 2.2.2 Coordinate space OPE
$C_{0}(z^{2}\mu^{2},d)=z^{4}(z^{2}\mu_{0}^{2})^{\frac{6-3d}{2}}\frac{2^{-d-5}\pi^{-\frac{3}{2}(d-1)}\Gamma\left(\frac{d}{2}-1\right)^{2}}{(9d^{2}-54d+80)\Gamma(5-d)\Gamma\left(\frac{d-1}{2}\right)\sin\pi d}\ ,$ (64)
$C_{1}(z^{2}\mu^{2},d)=-z^{4}(z^{2}\mu_{0}^{2})^{2-d}\frac{2^{-d-5}\pi^{\frac{3}{2}-d}\Gamma(d-4)}{\Gamma\left(4-\frac{d}{2}\right)\Gamma\left(\frac{d-1}{2}\right)\sin\left(\frac{\pi d}{2}\right)}\ ,$ (65)
$C_{2}(z^{2}\mu^{2},d)=z^{4}(z^{2}\mu_{0}^{2})^{1-\frac{d}{2}}\frac{\pi^{-\frac{d}{2}}}{128}\ \Gamma\left(\frac{d}{2}-3\right)\ ,$ (66)
@@
Given the above, one has
@@
which defines the counter-terms and renormalized operators. Defining the operators
@@
Here in the $L_{\mu}$, $m$ is replaced by the $\overline{MS}$ scheme $\mu$. In fact, by requiring the cancellation of $\frac{1}{\epsilon}$ poles, the three mixing terms from $C_{3}$
@@
3 Higher order behavior of OPE expansion in large-$N$ expansion: factorial enhancements and their cancellation
@@
For example, lets consider the coefficient function $C_{\rm I}$ for the identity operator that arises from $\phi(z)\times\phi(0)$. The $C_{\rm I}$ has an expansion
@@
where $C_{n,k}$ depends only on $L_{\mu}$. Here we focus on the terms $C_{n,k}$. These terms are nothing but the IR renormalized two-point functions in the massless perturbation theory at the order $n-k$ with $k$ mass insertions. Heuristically, one may argue that the total numbers of Feynman diagrams grow as $(n)!$, while being the $n$-th power in the small $z^{2}$ expansion introduces another factor of $\frac{1}{(n!)^{2}}$. Thus $C_{n,k}$ may still be supressed by $\frac{1}{n!}$, suggesting an infinite convergence radius for the power expansion. However, the fact that $C_{n,k}$ requires IR-renormalization, may introduces additional factorial growth, in a way similar to the famous IR-renormalon problem ’t Hooft (1979); Parisi (1979, 1978) for the logarithmic asymptotics in marginal theory.
To investigate this issue, one may again use the large-$N$ expansion trick. Consider the following $O(N)$ generalization of the scalar $\phi^{4}$ theory in 2D  [^2] In Mariño and Reis (2020), this model have been studied in the $O(N)$-broken false vacuum and a renormalon has been observed in the coupling-constant expansion of the free-energy. In our work, we are interested in the OPE expansion in the physical unbroken phase, and it remains to see if there are connections between the renormalon found in Mariño and Reis (2020) and the factorial enhancements observed in our work.
@@
Here $a=1,2,3,...N$ with summation convention implied, and we have use again the bare mass $m_{0}^{2}=m^{2}+\delta m^{2}-\frac{g}{2}I_{0}$. Here, the Wick-ordering is performed with respect to free Gaussian field with mass $m$. Notice that $\delta m^{2}={\cal O}\left(\frac{g}{N}\right)$ which we chose to cancel the one-loop self-energy correction at $\frac{1}{N}$. Equivalently, one may write in terms of the overall Wick-ordering $:(\Phi^{a}\Phi^{a})^{2}:$. Also notice that the normalization for the quartic term is different from that of the $N=1$. Unlike the $N=1$ case, for $N\geq 3$ it is normally conjectured in a way similar to the lattice $O(N)$ model Polyakov (1975) that there is no phase transition as one increases the coupling constant.
Here we consider the coefficient function for the identity operator in the OPE exapsion in the short distance limit $z^{2}g\ll 1,z^{2}m^{2}\ll 1$, for again the two point function $\langle\Phi^{a}(z)\Phi^{b}(0)\rangle\rightarrow\delta^{ab}C_{\rm I}(z)+....$. The IR structure is irrelevant for this limit. Clearly, one has the large $N$ expansion $C_{\rm I}=C_{\rm I,L}+\frac{1}{N}C_{\rm I,NL}+...$ To calculate $C_{\rm I,1}$, notice that we only need to consider a massless bubble chain and subtract out the IR sub-divergences to obtain the IR renormalized coefficient functions. To further demonstrates the factorial enhancement, and to make contact with previous investigations, in the following we perform the calculation in the momentum space expansion, but will comment on the features of the coordinate space expansion.

### v2

@@ 2.2.1 Momentum space OPE
Figure 4: The coefficient functions for various operators at the order ${\cal O}(z^{4})$. The dashed lines are massless propagators. The big black circles are insertion of the bare mass $-m_{0}^{2}$. The double lines denote the “amputated external soft fields” to keep track the operator content of the condensates. The tree-level coefficient $-\frac{3z^{4}}{24d(d+2)}$ for $\phi\partial^{4}\phi$ are omitted in the figure, but crucial for coordinate space expansion.
@@
$C_{0}\left(\frac{p^{2}}{\mu^{2}},d\right)=\mu_{0}^{4-2d}\int\frac{d^{d}kd^{d}l}{(2\pi)^{2d}}\frac{1}{k^{2}l^{2}(p-k-l)^{2}}$
$=\frac{1}{p^{2}}\left(\frac{\mu_{0}^{2}}{p^{2}}\right)^{2-d}\frac{8^{1-d}\pi^{\frac{3}{2}-d}\Gamma\left(\frac{d}{2}-1\right)^{2}}{\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{3d}{2}-3\right)\sin\pi d}\ ,$ (40)
$C_{1}\left(\frac{p^{2}}{\mu^{2}},d\right)=\mu_{0}^{2-d}\int\frac{d^{d}k}{(2\pi)^{d}}\frac{1}{k^{2}(p-k)^{2}}$
$=\frac{1}{p^{2}}\left(\frac{\mu_{0}^{2}}{p^{2}}\right)^{1-\frac{d}{2}}\frac{2^{3-2d}\pi^{\frac{1}{2}-\frac{d}{2}}\Gamma\left(2-\frac{d}{2}\right)\Gamma\left(\frac{d}{2}-1\right)}{\Gamma\left(\frac{d-1}{2}\right)}\ ,$ (41)
@@
Notice that $C_{0}$ and $C_{1}$ are just the massless self-energy diagrams with zero and two external field insertions without symmetry factors, see Fig. 4. One can show that
@@
which is exactly the expansion of the massive sunrise self-energy $\Sigma(p^{2})$ defined in Eq. (6). Clearly, this can be regarded as follows: $\Sigma(p^{2})$ at the power $\frac{1}{p^{2}}$ contains the following soft sub-regions: there are three ways for one of the three propagators in the sunrise to become soft and form the quadratic condensate $\langle\phi^{2}\rangle_{0}$. There are also three ways for two of them to form the quartic condensate $\langle\phi^{4}\rangle_{0}$. Combining with the region where all three propagators are hard, one indeed reproduces the expansion for $\Sigma(p^{2})$ at $\frac{1}{p^{2}}$.
We would like to proceed further to obtain the OPE for the two-point function. For this one must notice that the $\langle\phi^{4}\rangle_{0}$ insertion for the self-energy is not all of the $\langle\phi^{4}\rangle_{0}=3I_{0}^{2}$ contribution to the two-point function. The reason is that there is also a “reducible” term that should cancel with the $\frac{1}{\epsilon^{2}}$ for double $m_{0}^{2}$ insertion. Indeed, if one compute the last diagram in Fig. (4) as a correlation function between $\phi(-p)\phi(p)$ with four amputated zero-momentum scalar fields, then one can see the coefficient function for $\phi_{0}^{4}$ insertion should be $\frac{g^{2}}{4p^{2}}$. The contribution to the self-energy is $\frac{g^{2}}{6p^{2}}\times 3I_{0}^{2}$, while there is another $\frac{1}{12p^{2}}\times 3I_{0}^{2}$ that should cancel with mass insertions. Similarly, for the $\langle\phi^{2}\rangle_{0}$ insertion one should also include the $m_{0}^{2}\times g$ order diagram shown in Fig. 4. Therefore, the bare OPE for the two-point function reads
$S(p^{2})(p^{2})^{2}=\left(\frac{g^{2}}{6}C_{0}+\frac{m_{0}^{4}}{p^{2}}\right)+\left(\frac{g^{2}}{2}C_{1}+\frac{gm_{0}^{2}}{p^{2}}\right)\langle\phi^{2}\rangle_{0}+\left(\frac{g^{2}}{6p^{2}}+\frac{g^{2}}{12p^{2}}\right)\langle\phi^{4}\rangle_{0}\ .$ (45)
Then, one introduces the renormalized operators in the $\overline{MS}$ scheme as
$\langle\phi^{4}\rangle_{0}=\langle\phi^{4}\rangle_{r}+\frac{3}{2\pi\epsilon}\langle\phi^{2}\rangle_{r}+\frac{3}{(4\pi)^{2}\epsilon^{2}}\ .$ (47)
Notice that $\langle\phi^{4}\rangle_{r}=3\langle\phi^{2}\rangle_{r}^{2}$ at tree-level, despite possible $\frac{1}{\epsilon}\times\epsilon$ effects. The point is that $\langle\phi^{2}\rangle_{r}$ and $\langle\phi^{4}\rangle_{r}$ are finite in the $\epsilon\rightarrow 0$ limit, but in the $\epsilon\rightarrow 0$ limit they can still contain ${\cal O}(\epsilon)$ terms.
Given the above, by collecting coefficients for renormalized operators, one has
@@
In the $\epsilon\rightarrow 0$ limit, all of the above are finite. In fact, requiring that the $\epsilon\rightarrow 0$ limit exist uniquely determines all the $\frac{1}{\epsilon}$ poles in the mixing, once the tree-level mixing constant $Z_{10}^{(0)}=\frac{1}{4\pi\epsilon}$ for $\phi^{2}$ is determined from the coordinate space expansion at $z^{0}$ order, and the one-loop level mixing $m_{0}^{2}=m_{r}^{2}-\frac{g}{8\pi\epsilon}$ is fixed from the $z^{2}$ order consideration. In particular
@@
Notice that all the $m_{0}^{2}$ insertions in $C_{\rm I}$ again combine with the operator UV poles for $\phi^{2n}$ inserted along the same propagator to form the renormalized mass $m_{r}^{2}$. Anticipating that, in the coefficient function calculation, one can instead use $m_{r}^{2}$ and neglect many operator UV poles corresponding to “reducible” insertions of $(\phi^{2})^{n}$ along the same propagator. This is particularly convenient in the large $N$ calculation, since in such cases different contractions between the $(\phi^{2})^{n}$ usually have different large $N$ power-counting.
Given the OPE, to reproduce the full result, notice in the current case one has $m_{r}^{2}=m^{2}-\frac{g}{2}\langle\phi^{2}\rangle_{r}=m^{2}-\frac{g}{8\pi}\ln\frac{\mu^{2}}{m^{2}}+{\cal O}(\epsilon)$, and the tree-level operator condensates in the $\epsilon\rightarrow 0$ limit read
@@
Combining all above, one has the $\mu$-independent result $S_{2}(p^{2})(p^{2})^{2}=\frac{g^{2}}{32\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{m^{2}}+\frac{m^{4}}{p^{2}}=C_{{\rm I},r}^{(4)}(\mu^{2}=m^{2})$. This again reveals an important fact: if only tree-level condensates are required, then one can set $\mu=m$ to remove them and keep only the coefficient function for the identity operator. This is because at tree-level, $\langle\phi^{2n}\rangle$ always factorize into $\frac{(2n)!}{(n!)2^{n}}\langle\phi^{2}\rangle^{n}$, therefore after renormalization they always vanish at $\mu=m$. This is also true if some of the $\phi^{2}$s are replaced by $\phi(\partial^{2k}\phi)$, but only self-contractions are present, and only the “tadpole-type” high-order corrections to these operators are allowed in terms of the $\frac{1}{k^{2}+m_{r}^{2}}$ propagators. Later in Sec. (3) we will use this observation.
Here we should stress that if one only keeps the naive perturbative result $\frac{g^{2}}{6}C_{0}$ in $C_{{\rm I},r}$ without introducing additive counter-terms, then not only the $\frac{1}{\epsilon}$ poles no longer cancel, but the double log term in the finite part would also have the wrong coefficient $\frac{g^{2}}{16\pi^{2}p^{2}}\ln^{2}\frac{p^{2}}{\mu^{2}}$.
@@
It is clear now that all the $\frac{1}{\epsilon}$ poles cancel, the $\frac{1}{\epsilon^{2}}$ poles add to cancel the remaining $\frac{g^{2}}{2p^{2}}\frac{1}{(4\pi)^{2}\epsilon^{2}}$. The $\zeta_{2}$ term all cancel and the double-log terms combine to reproduce the full result for $C_{{\rm I},r}$. Therefore, in the DR calculation, it is important to keep the $d$ dependency of the bare coefficient functions to the end, and the $\epsilon\times\frac{1}{\epsilon}$ effects can be crucial.
@@
Here we note that the above is the general form of the large $p^{2}$ expansion at the next-to-next-leading order in $\frac{1}{p^{2}}$ for a generic $\phi_{2}^{4}$, where $\langle:\phi^{2}:\rangle$ and $\langle:\phi^{4}:\rangle$ can be non-perturbative. At small $g$, condensates for these two Wick-polynomials are of orders ${\cal O}(\frac{g^{2}}{m^{4}})$ and ${\cal O}\left(\frac{g}{m^{2}}\right)$. The double logarithmic dependency is always purely determined perturbatively.
@@ 2.2.2 Coordinate space OPE
$C_{0}(z^{2}\mu^{2},d)$ $=\mu_{0}^{6-3d}\int\frac{d^{d}pd^{d}kd^{d}l}{(2\pi)^{3d}}\frac{e^{ip\cdot z}}{(p^{2})^{2}k^{2}l^{2}(p-k-l)^{2}}$
$=z^{4}(z^{2}\mu_{0}^{2})^{\frac{6-3d}{2}}\frac{2^{-d-5}\pi^{-\frac{3}{2}(d-1)}\Gamma\left(\frac{d}{2}-1\right)^{2}}{(9d^{2}-54d+80)\Gamma(5-d)\Gamma\left(\frac{d-1}{2}\right)\sin\pi d}\ ,$ (63)
$C_{1}(z^{2}\mu^{2},d)$ $=\mu_{0}^{4-2d}\int\frac{d^{d}pd^{d}k}{(2\pi)^{2d}}\frac{e^{ip\cdot z}}{(p^{2})^{2}k^{2}(p-k)^{2}}$
$=-z^{4}(z^{2}\mu_{0}^{2})^{2-d}\frac{2^{-d-5}\pi^{\frac{3}{2}-d}\Gamma(d-4)}{\Gamma\left(4-\frac{d}{2}\right)\Gamma\left(\frac{d-1}{2}\right)\sin\left(\frac{\pi d}{2}\right)}\ ,$ (64)
$C_{2}(z^{2}\mu^{2},d)$ $=\mu_{0}^{2-d}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{e^{ip\cdot z}}{(p^{2})^{3}}$
$=z^{4}(z^{2}\mu_{0}^{2})^{1-\frac{d}{2}}\frac{\pi^{-\frac{d}{2}}}{128}\ \Gamma\left(\frac{d}{2}-3\right)\ ,$ (65)
@@
Notice that they are already for the two-point function rather than the self-energy. Given the above, one has
@@
which defines the UV poles and renormalized operators. Defining the operators
@@
Here in the $L_{\mu}$, $m$ is replaced by the $\overline{MS}$ scheme $\mu$. In fact, by requiring the cancellation of $\frac{1}{\epsilon}$ poles, the three series of UV poles for $O_{3}\rightarrow O_{i}$ mixing
@@
3 Higher order behavior of OPE expansion in a large-$N$ quartic $O(N)$ model: factorial enhancements and their cancellation
@@
For example, let’s consider the coefficient function $C_{\rm I}$ for the identity operator that arises from $\phi(z)\times\phi(0)$. The $C_{\rm I}$ has an expansion
@@
where $C_{n,k}$ depends only on $L_{\mu}$. Here we focus on the terms $C_{n,k}$. These terms are nothing but the IR renormalized two-point functions in the massless perturbation theory at the order $n-k$ with $k$ mass insertions. Heuristically, one may argue that the total numbers of Feynman diagrams grow as $(n)!$, while being the $n$-th power in the small $z^{2}$ expansion introduces another factor of $\frac{1}{(n!)^{2}}$. Thus $C_{n,k}$ may still be suppressed by $\frac{1}{n!}$, suggesting an infinite convergence radius for the power expansion. However, the fact that $C_{n,k}$ requires IR-renormalization, may introduce additional factorial growth, in a way similar to the famous IR-renormalon problem ’t Hooft (1979); Parisi (1979, 1978) for the logarithmic asymptotics in marginal theory.
To investigate this issue, one may again use the large-$N$ expansion trick. Consider the following $O(N)$ generalization of the scalar $\phi^{4}$ theory in 2D  [^2] In Mariño and Reis (2020), this model has been studied in the $O(N)$-broken false vacuum and a renormalon has been observed in the coupling-constant expansion of the free-energy. In our work, we are interested in the OPE expansion in the physical unbroken phase, and it remains to see if there are connections between the renormalon found in Mariño and Reis (2020) and the factorial enhancements observed in our work.
@@
Here $a=1,2,3,...N$ with summation convention implied, and we have use again the bare mass $m_{0}^{2}=m^{2}+\delta m^{2}-\frac{g}{2}I_{0}$. Here, the Wick-ordering is performed with respect to the free Gaussian field with mass $m$. Notice that $\delta m^{2}={\cal O}\left(\frac{g}{N}\right)$ which we chose to cancel the one-loop self-energy correction at $\frac{1}{N}$. Equivalently, one may write in terms of the overall Wick-ordering $:(\Phi^{a}\Phi^{a})^{2}:$. Also notice that the normalization for the quartic term is different from that of the $N=1$. Unlike the $N=1$ case, for $N\geq 3$ it is normally conjectured in a way similar to the lattice $O(N)$ model Polyakov (1975) that there is no phase transition as one increases the coupling constant. To simplify notation, from now on we use
$\Phi^{2}\equiv\sum_{a=1}^{N}\Phi^{a}\Phi^{a}\ ,\ \Phi(i\partial)^{n}\Phi\equiv\sum_{a=1}^{N}\Phi^{a}(i\partial)^{n}\Phi^{a}\ ,$ (88)
to denote $O(N)$-contracted quadratic operators.
Here we consider the coefficient function for the identity operator in the OPE exapsion in the short distance limit $z^{2}g\ll 1,z^{2}m^{2}\ll 1$, for again the two point function $\langle\Phi^{a}(z)\Phi^{b}(0)\rangle\rightarrow\delta^{ab}C_{\rm I}(z)+....$. The IR structure is irrelevant for this limit. Clearly, one has the large $N$ expansion $C_{\rm I}=C_{\rm I,L}+\frac{1}{N}C_{\rm I,NL}+...$ To calculate $C_{\rm I,{\rm NL}}$, notice that we only need to consider a massless bubble chain and subtract out the IR sub-divergences to obtain the IR renormalized coefficient functions. To further amplify the factorial enhancements, and to make contact with previous investigations, in the following we perform the calculation in the momentum space expansion, but will comment on the features of the coordinate space expansion.